solve the system of linear equations and check the solution
3y=11-2x
3x=y-11

Answer by macston(5194)

.
3x=y-11
3x+11=y
.
3y=11-2x
3(3x+11)=11-2x
9x+33=11-2x
11x=-22
x=-2
.
3x=y-11
3(-2)=y-11
-6=y-11
5=y
.
CHECK:
3y=11-2x
3(5)=11-2(-2)
15=11+4
15=15